Stuck on this problem, it would be appreciated if steps were shown. thanks!
A rectangle has a perimeter of 46 inches. The width of the rectangle is 7 inches shorter than twice its length. If a square's side is as long as the width of this rectangle, find the perimeter of the square.

Answer by mananth(16949)

Let length = x
width = 2x-7
Perimeter = 2(L+W)
Perimeter = 2(x+2x-7)
46= 2(3x-7)
46=6x-14
+14
60 = 6x
/6
x= 10 inches the length
width = 2x-7 = 20-7 = 13
The perimeter of square = 4*13 = 52 inches